# Artificial barriers for stochastic differential equations and for construction of boundary-preserving schemes

### v2

@@ 5. Artificial Barriers Euler–Maruyama
$\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t)$ $=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k})+\int_{t_{k}}^{t}f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}s+\int_{t_{k}}^{t}g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}B(s)$
$=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)+\int_{t_{k}}^{t}f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}s+\int_{t_{k}}^{t}g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}B(s)$
$=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)+f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))(t-t_{k})+g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))(B(t)-B(t_{k}))$
for $t\in[t_{k},t_{k+1})$. Moving $\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)$ to the left hand side we thus have
${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{k})-\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t)\right|^{p}\right]$ $\leq C(p){\mathbb{E}}\left[\sup_{k=0,\ldots,M-1}\sup_{t\in[t_{k},t_{k+1})}\left|f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{k-1}))(t-t_{k})\right|^{p}\right]$
$+C(p){\mathbb{E}}\left[\sup_{k=0,\ldots,M-1}\sup_{t\in[t_{k},t_{k+1})}\left|g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{k-1}))(B(t)-B(t_{k}))\right|^{p}\right]$
The first term on the right hand side can be bounded by $C(p,T,L_{f},L_{g},a,x_{0})\Delta t^{p}$ if $b=\infty$ and by $C(p,L_{f},a,b)\Delta t^{p}$ if $|b|<\infty$ by linear growth of $f$ on $D$ in (3) and the moment bounds on $\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}$ in Proposition 24. For the second term, we first use Cauchy-Schwarz inequality to separate the term involving $g$ and the term involving $B(t)-B(t_{k})$ as such:
${\mathbb{E}}\left[\sup_{k=0,\ldots,M-1}\sup_{t\in[t_{k},t_{k+1})}\left|g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{k-1}))(B(t)-B(t_{k}))\right|^{p}\right]$ $\leq\left({\mathbb{E}}\left[\sup_{k=0,\ldots,M-1}\sup_{t\in[t_{k},t_{k+1})}\left|g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{k-1}))\right|^{2p}\right]\right)^{1/2}$
×$\times\left({\mathbb{E}}\left[\sup_{k=0,\ldots,M-1}\sup_{t\in[t_{k},t_{k+1})}\left|(B(t)-B(t_{k}))\right|^{2p}\right]\right)^{1/2}$
@@
(33) $\left({\mathbb{E}}\left[\sup_{k=0,\ldots,M-1}\sup_{t\in[t_{k},t_{k+1})}\left|(B(t)-B(t_{k}))\right|^{2p}\right]\right)^{1/2}\leq C(p,T)\Delta t^{p/2-\epsilon},$
@@
(34) ${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{k})-\tilde{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t)\right|^{p}\right]\leq C(p,T,L_{f},L_{g},a,b,x_{0})\Delta t^{p/2-\epsilon},$
@@
$\leq C(p){\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\tilde{w}_{N,\Delta t}(t)-w_{N}(t)\right|^{p}\right].$
@@
$\sup_{t\in[0,T]}|\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t)-X(t)|\leq\eta_{\epsilon}\Delta t^{1/2-\epsilon}$
almost surely. Moreover, $\eta_{\epsilon}$ is independent of $N\in\mathbb{N}$.
@@ 6. Artificial Barriers Euler–Peano
Next, we combine the method of artificial barriers developed in Section 4 with a modification of the classical Euler–Peano (EP) scheme for RSDEs applied to the RSDE with modified coefficients in (23). We refer to the resulting scheme as the Artificial Barriers Euler–Peano scheme and we abbreviate it as ABEP. We first provide the definition of the scheme and then we verify that the ABEP scheme satisfy Assumption 3 in Section 4. By Theorem 22, we can then conclude that the ABEP scheme is boundary-preserving and converges to the solution $X$ of the considered SDE (1).
@@
We would like to make an important remark before proceeding to defining the ABEP scheme. There are explicit expressions for the EP and ABEP schemes (see equation (37) for the ABEP scheme). However, as the supremum and infimum in Definition 3 and Definition 5 are over infinitely many time points, the EP and ABEP schemes are not (currently) implementable. In other words, as of now, the ABEP scheme is of more theoretical than practical use. We include the ABEP scheme as it may be used to construct other implementable schemes (for example, a modified version of the Euler scheme for RSDEs studied in [24]).
We now provide a step-by-step construction of the ABEP scheme. We refer the reader not familiar with RSDEs to equation (37) for a representation of the ABEP scheme without any mentioning of RSDEs. Recall that the classical EM scheme for the SDE (2) is given by
$X^{\textrm{EM}}_{\Delta t}(t_{k+1})=X^{\textrm{EM}}_{\Delta t}(t_{k})+f(X^{\textrm{EM}}_{\Delta t}(t_{k}))\Delta t+g(X^{\textrm{EM}}_{\Delta t}(t_{k}))\Delta B_{k}$
and, for the purpose of defining the ABEP scheme, we extend the above to continuous time by requiring that $X^{\textrm{EM}}_{\Delta t}(t)$ solves the following SDE with piecewise constant coefficients
$\mathop{}\!\mathrm{d}X^{\textrm{EM}}_{\Delta t}(t)=f(X^{\textrm{EM}}_{\Delta t}(\ell_{\Delta t}(t)))\mathop{}\!\mathrm{d}t+g(X^{\textrm{EM}}_{\Delta t}(\ell_{\Delta t}(t)))\mathop{}\!\mathrm{d}B(t),$
where we recall that $\ell_{\Delta t}(t)=t_{k}$ for $t\in[t_{k},t_{k+1})$, with initial value $X^{\textrm{EM}}_{\Delta t}(0)=x_{0}$. We now add a lower barrier process $\hat{L}_{N,\Delta t}^{\textrm{ABEP}}$ and an upper barrier process $\hat{U}_{N,\Delta t}^{\textrm{ABEP}}$, in the sense of Section 3.2, to force $X^{\textrm{EM}}_{\Delta t}$ to stay in $\operatorname{cl}D_{N}$:
$\mathop{}\!\mathrm{d}\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(t)=f(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(t)))\mathop{}\!\mathrm{d}t+g(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(t)))\mathop{}\!\mathrm{d}B(t)-\mathop{}\!\mathrm{d}\hat{U}_{N,\Delta t}^{\textrm{ABEP}}(t)+\mathop{}\!\mathrm{d}\hat{L}_{N,\Delta t}^{\textrm{ABEP}}(t)$
@@
(35) $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(t)=x_{0}+\int_{0}^{t}f(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(s)))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(s)))\mathop{}\!\mathrm{d}B(s)-\hat{U}_{N,\Delta t}^{\textrm{ABEP}}(t)+\hat{L}_{N,\Delta t}^{\textrm{ABEP}}(t).$
@@
(36) $\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)=x_{0}+\int_{0}^{t}f(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(s)))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(s)))\mathop{}\!\mathrm{d}B(s)$
to obtain an associated triple $(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t},\hat{w}_{N,\Delta t}^{\textrm{ABEP}},\hat{L}_{N,\Delta t}^{\textrm{ABEP}}-\hat{U}_{N,\Delta t}^{\textrm{ABEP}})$ in the sense of Section 3.2. Then, in particular, we have that $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}=\Gamma_{N}(\hat{w}_{N,\Delta t}^{\textrm{ABEP}})$. Note that (36) on $[\ell_{\Delta t}(s),\ell_{\Delta t}(s)+\Delta t]$ is the classical EM scheme starting from the value $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(s))$. Moreover, using the explicit expressions for the Skorokhod map in Definition 5, we have the following representation for $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}$ in the two-finite-boundaries case ($|b|<\infty$)
(37) $\begin{split}\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(t)&=a+\frac{1}{N}+\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)+\sup_{s\in[0,t]}\max(0,-\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(s))\\ &-\sup_{s\in[0,t]}\left(\min\left(\max\left(0,\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(s)+\sup_{r\in[0,s]}\max\left(0,-\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(r)\right)-\left(b-a-\frac{2}{N},\right)\right),\right.\right.\\ &\left.\left.\inf_{u\in[s,t]}\left(\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(u)+\sup_{v\in[0,u]}\max\left(0,-\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(v)\right)\right)\right)\right)\end{split}$
@@
The following proposition provides moment bounds and boundary-preservation of the ABEP scheme. Since the ABEP scheme is defined in (35) as the solution to a reflected SDE, the boundary-preserving property of the ABEP scheme follows directly.
@@
${\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\hat{\xi}_{N,\Delta t}^{\textrm{ABEP}}(t)\right|^{p}\right]={\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\Gamma_{N}\left(\hat{w}_{N,\Delta t}^{\textrm{ABEP}}\right)(t)\right|^{p}\right]\leq C(p){\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)\right|^{p}\right].$
@@
$I_{1}\leq C(p,T,L_{f},a)\int_{0}^{r}\sup_{t\in[0,s]}|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)|^{p}\mathop{}\!\mathrm{d}s$
@@
${\mathbb{E}}\left[I_{2}\right]\leq C(p,T,L_{g},a)\int_{0}^{r}{\mathbb{E}}\left[\sup_{t\in[0,s]}|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)|^{p}\right]\mathop{}\!\mathrm{d}s.$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)|^{p}\right]\leq C(p,T,L_{f},L_{g},a)\left(|x_{0}|^{p}+\int_{0}^{r}{\mathbb{E}}\left[\sup_{t\in[0,s]}|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)|^{p}\right]\mathop{}\!\mathrm{d}s\right)$
@@
We now state and prove the convergence theorem of this section; namely, the $L^{p}(\Omega)$-convergence, for every $p\geq 2$, of $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}$ to the solution $X$ of the considered SDE (1). As for Theorem 25, the proof of Theorem 28 verifies that Assumption 3 is fulfilled and the result then follows from Theorem 22. Note that convergence of the Euler–Peano scheme is known (see, for example, [40]), we include it since the coefficient functions of the RSDE for $\xi_{N}$ are not the same as in the definition of the modified Euler–Peano scheme $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}$.
@@
for every $p\geq 2$, where the constant $C(p,T,L_{f},L_{g},a,b,x_{0})$ is independent of $N\in\mathbb{N}$ and $\Delta t$. If $b=\infty$, then the constant is independent of $b$.
@@
(38) ${\mathbb{E}}\left[\sup_{t\in[0,T]}|\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(t)-\xi_{N}(t)|^{p}\right]\leq C(p,T,L_{f},L_{g})\left(N^{-p}+\Delta t^{p/2}\right),$
for some constant independent of $N\in\mathbb{N}$ and independent of $\Delta t>0$.
@@
(39) $\begin{split}\sup_{t\in[0,r]}&|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)-w_{N}(t)|^{p}\leq\\ &C(p)\sup_{t\in[0,r]}\left|\int_{0}^{t}f(\Gamma_{N}(\hat{w}_{N,\Delta t}^{\textrm{ABEP}})(\ell_{\Delta t}(s)))-f_{N}(\Gamma_{N}(w_{N})(s))\mathop{}\!\mathrm{d}s\right|^{p}\\ +&C(p)\sup_{t\in[0,r]}\left|\int_{0}^{t}g(\Gamma_{N}(\hat{w}_{N,\Delta t}^{\textrm{ABEP}})(\ell_{\Delta t}(s)))-g_{N}(\Gamma_{N}(w_{N})(s))\mathop{}\!\mathrm{d}B(s)\right|^{p}\\ =&C(p)(J_{1}+J_{2}).\end{split}$
${\mathbb{E}}\left[J_{1}\right]$ and ${\mathbb{E}}\left[J_{2}\right]$ are estimated in the same way, the difference is that we use Lemma 15 for ${\mathbb{E}}\left[J_{1}\right]$ and we use Lemma 17 for ${\mathbb{E}}\left[J_{2}\right]$ We provide the argument for ${\mathbb{E}}\left[J_{2}\right]$, as the term involving the noise is (usually) the difficult part. Using (remarks to) Lemma 17, the Lipschitz estimate for $\Gamma_{N}$ in (14) and moment bounds for $\hat{\xi}_{N,\Delta t}^{\textrm{ABEP}}=\Gamma_{N}(\hat{w}_{N,\Delta t}^{\textrm{ABEP}})$ in Theorem 27, we can estimate
@@
$+C(p,T,L_{g},a,b)N^{-2}\left(1+{\mathbb{E}}\left[\sup_{t\in[0,T]}|\Gamma_{N}(\hat{w}_{N,\Delta t}^{\textrm{ABEP}})(t)|^{p}\right]\right)$

### v3

@@ 5. Artificial Barriers Euler–Maruyama
$\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t)$ $=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t_{m})+\int_{t_{m}}^{t}f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}s+\int_{t_{m}}^{t}g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}B(s)$
$=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)+\int_{t_{m}}^{t}f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}s+\int_{t_{m}}^{t}g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(s))\mathop{}\!\mathrm{d}B(s)$
$=\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)+f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{m}))(t-t_{m})+g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{m}))(B(t)-B(t_{m}))$
for $t\in[t_{m},t_{m+1})$. Moving $\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)$ to the left hand side we thus have
${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\hat{w}_{N,\Delta t}^{\textrm{ABEM}}(t)-\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t)\right|^{p}\right]$ $\leq C(p){\mathbb{E}}\left[\sup_{m=0,\ldots,M-1}\sup_{t\in[t_{m},t_{m+1})}\left|f(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{m}))(t-t_{m})\right|^{p}\right]$
$+C(p){\mathbb{E}}\left[\sup_{m=0,\ldots,M-1}\sup_{t\in[t_{m},t_{m+1})}\left|g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{m}))(B(t)-B(t_{m}))\right|^{p}\right]$
The first term on the right hand side can be bounded by $C(p,T,L_{f},L_{g},a,x_{0})\Delta t^{p}$ if $b=\infty$ and by $C(p,L_{f},a,b)\Delta t^{p}$ if $|b|<\infty$ by linear growth of $f$ on $D$ in (3) and the moment bounds on $\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}$ in Proposition 24. For the second term, we first use Cauchy-Schwarz inequality to separate the term involving $g$ and the term involving $B(t)-B(t_{m})$ as such:
${\mathbb{E}}\left[\sup_{m=0,\ldots,M-1}\sup_{t\in[t_{m},t_{m+1})}\left|g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{m-1}))(B(t)-B(t_{m}))\right|^{p}\right]$ $\leq\left({\mathbb{E}}\left[\sup_{m=0,\ldots,M-1}\sup_{t\in[t_{m},t_{m+1})}\left|g(\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{m}))\right|^{2p}\right]\right)^{1/2}$
$\times\left({\mathbb{E}}\left[\sup_{m=0,\ldots,M-1}\sup_{t\in[t_{m},t_{m+1})}\left|(B(t)-B(t_{m}))\right|^{2p}\right]\right)^{1/2}$
@@
(34) $\left({\mathbb{E}}\left[\sup_{m=0,\ldots,M-1}\sup_{t\in[t_{m},t_{m+1})}\left|(B(t)-B(t_{m}))\right|^{2p}\right]\right)^{1/2}\leq C(p,T)\Delta t^{p/2-\epsilon},$
@@
(35) ${\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\hat{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t_{m})-\tilde{\xi}_{N,\Delta t}^{\textrm{ABEM}}(t)\right|^{p}\right]\leq C(p,T,L_{f},L_{g},a,b,x_{0})\Delta t^{p/2-\epsilon},$
@@
$\leq C(p){\mathbb{E}}\left[\sup_{t\in[0,r]}\left|\tilde{w}_{N,\Delta t}^{\textrm{ABEM}}(t)-w_{N}(t)\right|^{p}\right].$
@@
Suppose Assumptions 1 and 2 are satisfied and let $N\in\mathbb{N}$ and $\Delta t=N^{-2}$. Then there exists, for every $\epsilon>0$, a random variable $\eta_{\epsilon}$, with ${\mathbb{E}}\left[\eta_{\epsilon}^{p}\right]<\infty$ for every $p\geq 1$, such that
$\sup_{t\in[0,T]}|\hat{\xi}^{\textrm{ABEM}}_{N,\Delta t}(t)-X(t)|\leq\eta_{\epsilon}\Delta t^{1/2-\epsilon},$
almost surely. Moreover, $\eta_{\epsilon}$ is independent of $N\in\mathbb{N}$ and $\Delta t$.
@@ 6. Artificial Barriers Euler–Peano
In this section, we combine the method of artificial barriers developed in Section 4 with a modification of the classical Euler–Peano (EP) scheme for RSDEs applied to the RSDE with modified coefficients in (24). We refer to the resulting scheme as the Artificial Barriers Euler–Peano scheme and we abbreviate it as ABEP. We first provide the definition of the scheme and then we verify that the ABEP scheme satisfy Assumption 3 in Section 4. By Theorem 22, we can then conclude that the ABEP scheme is boundary-preserving and converges to the solution $X$ of the considered SDE (1).
@@
We would like to make an important remark before proceeding to the definition of the ABEP scheme. There are explicit expressions for the EP and ABEP schemes (see equation (38) for the ABEP scheme). However, as the supremum and infimum in Definition 3 and Definition 5 are over infinitely many time points, the EP and ABEP schemes are not known to be implementable. In other words, as of now, the ABEP scheme is of more theoretical than practical use. We include the ABEP scheme as it may be used to construct other implementable schemes (for example, a modified version of the Euler scheme for RSDEs studied in [29]).
We now provide a step-by-step construction of the ABEP scheme. We refer the reader not familiar with RSDEs to equation (38) for a representation of the ABEP scheme without any mentioning of RSDEs. Recall that the classical EM scheme for the SDE in (2) is given by
$X^{\textrm{EM}}_{\Delta t}(t_{m+1})=X^{\textrm{EM}}_{\Delta t}(t_{m})+f(X^{\textrm{EM}}_{\Delta t}(t_{m}))\Delta t+g(X^{\textrm{EM}}_{\Delta t}(t_{m}))\Delta B_{m},$
for $m=0,\ldots,M-1$, and, for the purpose of defining the ABEP scheme, we extend the above to continuous time by requiring that $X^{\textrm{EM}}_{\Delta t}(t)$ solves the following SDE with piecewise constant coefficients
$\mathop{}\!\mathrm{d}X^{\textrm{EM}}_{\Delta t}(t)=f(X^{\textrm{EM}}_{\Delta t}(\ell_{\Delta t}(t)))\mathop{}\!\mathrm{d}t+g(X^{\textrm{EM}}_{\Delta t}(\ell_{\Delta t}(t)))\mathop{}\!\mathrm{d}B(t),\ t\in[0,T],$
where we recall that $\ell_{\Delta t}(t)=t_{m}$ for $t\in[t_{m},t_{m+1})$, with initial value $X^{\textrm{EM}}_{\Delta t}(0)=x_{0}$. We now add a lower barrier process $\hat{L}_{N,\Delta t}^{\textrm{ABEP}}$ and an upper barrier process $\hat{U}_{N,\Delta t}^{\textrm{ABEP}}$, in the sense of Section 3.2, to force $X^{\textrm{EM}}_{\Delta t}$ to stay in $\operatorname{cl}D_{N}$:
$\mathop{}\!\mathrm{d}\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(t)=f(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(t)))\mathop{}\!\mathrm{d}t+g(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(t)))\mathop{}\!\mathrm{d}B(t)-\mathop{}\!\mathrm{d}\hat{U}_{N,\Delta t}^{\textrm{ABEP}}(t)+\mathop{}\!\mathrm{d}\hat{L}_{N,\Delta t}^{\textrm{ABEP}}(t),\ t\in[0,T],$
@@
(36) $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(t)=x_{0}+\int_{0}^{t}f(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(s)))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(s)))\mathop{}\!\mathrm{d}B(s)-\hat{U}_{N,\Delta t}^{\textrm{ABEP}}(t)+\hat{L}_{N,\Delta t}^{\textrm{ABEP}}(t),\ t\in[0,T].$
@@
(37) $\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)=x_{0}+\int_{0}^{t}f(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(s)))\mathop{}\!\mathrm{d}s+\int_{0}^{t}g(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(s)))\mathop{}\!\mathrm{d}B(s),\ t\in[0,T],$
to obtain an associated triple $(\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t},\hat{w}_{N,\Delta t}^{\textrm{ABEP}},\hat{L}_{N,\Delta t}^{\textrm{ABEP}}-\hat{U}_{N,\Delta t}^{\textrm{ABEP}})$ in the sense of Section 3.2. Then, in particular, we have that $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}=\Gamma_{N}(\hat{w}_{N,\Delta t}^{\textrm{ABEP}})$. Note that (37) on $\ell_{\Delta t}(s),\ell_{\Delta t}(s)+\Delta t]$ is the classical EM scheme starting from the value $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(\ell_{\Delta t}(s))$. Moreover, using the explicit expressions for the Skorokhod map in Definition 5, we have the following representation for $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}$ in the two-finite-boundaries case ($|b|<\infty$)
(38) $\begin{split}\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(t)&=a+\frac{1}{N}+\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)+\sup_{s\in[0,t]}\max(0,-\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(s))\\ &-\sup_{s\in[0,t]}\left(\min\left(\max\left(0,\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(s)+\sup_{r\in[0,s]}\max\left(0,-\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(r)\right)-\left(b-a-\frac{2}{N},\right)\right),\right.\right.\\ &\left.\left.\inf_{u\in[s,t]}\left(\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(u)+\sup_{v\in[0,u]}\max\left(0,-\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(v)\right)\right)\right)\right),\ t\in[0,T].\end{split}$
@@
The following proposition provides moment bounds and boundary-preservation of the ABEP scheme. Since the ABEP scheme is defined in (36) as the solution to an RSDE, the boundary-preserving property of the ABEP scheme follows directly.
@@
${\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\hat{\xi}_{N,\Delta t}^{\textrm{ABEP}}(t)\right|^{p}\right]={\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\Gamma_{N}\left(\hat{w}_{N,\Delta t}^{\textrm{ABEP}}\right)(t)\right|^{p}\right]\leq C(p)\left({\mathbb{E}}\left[\sup_{t\in[0,T]}\left|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)\right|^{p}\right]+|a|^{p}+1\right).$
@@
$I_{1}\leq C(p,T,L_{f})\left(\int_{0}^{r}\sup_{t\in[0,s]}|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)|^{p}\mathop{}\!\mathrm{d}s+|a|^{p}+1\right)$
@@
${\mathbb{E}}\left[I_{2}\right]\leq C(p,T,L_{g})\left(\int_{0}^{r}{\mathbb{E}}\left[\sup_{t\in[0,s]}|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)|^{p}\right]\mathop{}\!\mathrm{d}s+|a|^{p}+1\right).$
@@
${\mathbb{E}}\left[\sup_{t\in[0,r]}|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)|^{p}\right]\leq C(p,T,L_{f},L_{g})\left(|x_{0}|^{p}+\int_{0}^{r}{\mathbb{E}}\left[\sup_{t\in[0,s]}|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)|^{p}\right]\mathop{}\!\mathrm{d}s+|a|^{p}+1\right)$
@@
We now state and prove the convergence theorem of this section; namely, the $L^{p}(\Omega)$-convergence, for every $p\geq 2$, of $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}$ to the solution $X$ of the considered SDE (1). As for Theorem 25, the proof of Theorem 28 verifies that Assumption 3 is fulfilled and the result then follows from Theorem 22. Note that convergence of the EP scheme is known (see, for example, [48]), we include it since, in contrast to the EP scheme, the coefficient functions of the RSDE for $\xi_{N}$ are not the same as in the definition of the modified EP scheme $\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}$.
@@
for every $p\geq 2$, where the constant is independent of $N\in\mathbb{N}$ and $\Delta t$. If $b=\infty$, then the constant is independent of $b$.
@@
(39) ${\mathbb{E}}\left[\sup_{t\in[0,T]}|\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(t)-\xi_{N}(t)|^{p}\right]\leq C(p,T,L_{f},L_{g},a,b,x_{0})\left(N^{-p}+\Delta t^{p/2}\right),$
for some constant independent of $N\in\mathbb{N}$ and independent of $\Delta t>0$, and independent of $b$ if $b=\infty$.
@@
(40) $\begin{split}\sup_{t\in[0,r]}&|\hat{w}_{N,\Delta t}^{\textrm{ABEP}}(t)-w_{N}(t)|^{p}\\ &\leq C(p)\sup_{t\in[0,r]}\left|\int_{0}^{t}f(\Gamma_{N}(\hat{w}_{N,\Delta t}^{\textrm{ABEP}})(\ell_{\Delta t}(s)))-f_{N}(\Gamma_{N}(w_{N})(s))\mathop{}\!\mathrm{d}s\right|^{p}\\ &+C(p)\sup_{t\in[0,r]}\left|\int_{0}^{t}g(\Gamma_{N}(\hat{w}_{N,\Delta t}^{\textrm{ABEP}})(\ell_{\Delta t}(s)))-g_{N}(\Gamma_{N}(w_{N})(s))\mathop{}\!\mathrm{d}B(s)\right|^{p}\\ &=C(p)(J_{1}+J_{2}).\end{split}$
${\mathbb{E}}\left[J_{1}\right]$ and ${\mathbb{E}}\left[J_{2}\right]$ are estimated in the same way, the difference is that we use Lemma 15 for ${\mathbb{E}}\left[J_{1}\right]$ and we use Lemma 17 for ${\mathbb{E}}\left[J_{2}\right]$ . We provide the argument for ${\mathbb{E}}\left[J_{2}\right]$, as the term involving the noise is (usually) the difficult part. Using (remarks to) Lemma 17, the Lipschitz estimate for $\Gamma_{N}$ in (14), and moment bounds for $\hat{\xi}_{N,\Delta t}^{\textrm{ABEP}}=\Gamma_{N}(\hat{w}_{N,\Delta t}^{\textrm{ABEP}})$ in Theorem 27, we can estimate
@@
$+C(p,T,L_{g},a,b)N^{-p}\left(1+{\mathbb{E}}\left[\sup_{t\in[0,T]}|\Gamma_{N}(\hat{w}_{N,\Delta t}^{\textrm{ABEP}})(t)|^{p}\right]\right)$